### v1

# On the smallest eigenvalues of $3$-colorable graphs

Abstract
We prove that the set of the smallest eigenvalues attained by $3$-colorable graphs is dense in $(-\infty,-\lambda^{*})$, where $\lambda^{*}=\rho^{1/2}+\rho^{-1/2}\approx 2.01980$ and $\rho$ is the positive real root of $x^{3}=x+1$. As a consequence, in the context of spherical two-distance sets, our result precludes any further refinement of the forbidden-subgraph method through the chromatic number of signed graphs.
1 Introduction
Motivated by the problem of estimating the maximum size of spherical two-distance sets with prescribed inner products, Jiang and Polyanskii [6] studied certain families of signed graphs within the class $\mathcal{G}_{p}$ of $p$-colorable signed graphs. Here, a signed graph is a graph $G$ whose edges are each labeled by $+$ or --, and a signed graph $G$ is $p$-colorable if there exists a $p$-coloring of $V(G)$ such that the endpoints of every negative edge receive identical colors, and the endpoints of every positive edge receive different colors.
Remark.
Our coloring notion is the sign-reversed version of Cartwright–Harary [2], where negative edges must join vertices of different colors and positive edges of identical colors. We choose the current definition because we need to discuss the (unsigned) graph coloring as well, and the current definition is a generalization of that. By flipping the signs of all edges, one can go back and forth between the two different definitions, as we will do when restating various results from [6, 8].
We now introduce the key families of signed graphs. Given $\lambda\in\mathbb{R}$, let $\mathcal{G}_{p}(\lambda)$ denote the family of $p$-colorable signed graphs whose smallest eigenvalues are at least $-\lambda$. The smallest eigenvalue of a signed graph $G$, denoted by $\lambda_{1}(G)$, is the smallest eigenvalue of its signed adjacency matrix.
The Cauchy interlacing theorem implies that $\mathcal{G}_{p}(\lambda)$ is closed under taking induced subgraphs. Jiang and Polyanskii asked in [6] whether it is possible to define each of these families by a finite set of forbidden subgraphs.
Definition 1.
Given a class $\mathcal{G}$ of signed graphs, and a family $\mathcal{H}\subseteq\mathcal{G}$ that is closed under taking induced subgraphs, a set $\mathcal{F}\subseteq\mathcal{G}$ of signed graphs is a forbidden subgraph characterization of $\mathcal{H}$ within $\mathcal{G}$ if $\mathcal{H}$ consists exactly of signed graphs in $\mathcal{G}$ that do not contain any member of $\mathcal{F}$ as an induced subgraph.
Problem 2 (Problem 5.3 of Jiang and Polyanskii [6]).
For every $p\in\mathbb{N}^{+}$, determine the set of $\lambda\in\mathbb{R}$ for which $\mathcal{G}_{p}(\lambda)$ has a finite forbidden subgraph characterization within $\mathcal{G}_{p}$.
For $p\in\left\{1,2\right\}$, 2 is essentially resolved by [5, Theorem 1]. In this paper, we completely resolve the rest of 2 as follows.
Theorem 3.
For every $p\geq 3$, within the class $\mathcal{G}_{p}$ of $p$-colorable signed graphs, the family $\mathcal{G}_{p}(\lambda)$ of signed graphs with smallest eigenvalues at least $-\lambda$ has a finite forbidden subgraph characterization if and only if $\lambda<\lambda^{*}$, where
$\lambda^{*}:=\rho^{1/2}+\rho^{-1/2}\approx 2.01980,$
and $\rho$ is the unique real root of $x^{3}=1+x$.
The key ingredient is the following construction of $3$-colorable graphs.
Theorem 4.
For every $\lambda>\lambda^{*}$ and every $\varepsilon>0$, there exists a $3$-colorable graph $G$ such that $\lambda_{1}(G)\in(-\lambda-\varepsilon,-\lambda)$.
The rest of the paper is organized as follows. Section 2 derives Theorem 3 from Theorem 4. Section 3 proves Theorem 4, deferring two technical lemmas to Section 4. Finally, Section 5 revisits the spherical two-distance set motivation.
2 Forbidden subgraph characterization
We first briefly explain why 2 for $p\in\left\{1,2\right\}$ is essentially resolved before.
Theorem 5 (Theorem 1 of Jiang and Polyanskii [5]).
For every integer $m\geq 2$, let $\beta_{m}$ be the largest root of $x^{m+1}=1+x+\dots+x^{m-1}$, and let $\alpha_{m}:=\beta_{m}^{1/2}+\beta_{m}^{-1/2}$. The family of graphs with spectral radius at most $\lambda$ has a finite forbidden subgraph characterization (within the class of graphs) if and only if $\lambda<\lambda^{\prime}$ and $\lambda\notin\left\{\alpha_{2},\alpha_{3},\dots\right\}$, where
$\lambda^{\prime}:=\lim_{m\to\infty}\alpha_{m}=\varphi^{1/2}+\varphi^{-1/2}=\sqrt{2+\sqrt{5}}\approx 2.05817,$
and $\varphi$ is the golden ratio $(1+\sqrt{5})/2$. ∎
To reduce the case where $p=2$ to $p=1$, we rely on a useful tool in spectral graph theory for signed graphs — two signed graphs are switching equivalent if one graph can be obtained from the other by reversing all the edges in a cut-set. An important feature of switching equivalence is that the switching equivalent signed graphs all have the same spectrum.
Corollary 6.
For $p\in\left\{1,2\right\}$, within the class $\mathcal{G}_{p}$ of $p$-colorable signed graphs, the family $\mathcal{G}_{p}(\lambda)$ of signed graphs with smallest eigenvalues at least $-\lambda$ has a finite forbidden subgraph characterization if and only if $\lambda<\lambda^{\prime}$ and $\lambda\not\in\left\{\alpha_{2},\alpha_{3},\dots\right\}$.
Proof.
Notice that for every signed graph $G$, it is $1$-colorable if and only if $-G$ is unsigned, and $\lambda_{1}(G)\geq-\lambda$ if and only if the largest eigenvalue of $-G$ is at most $\lambda$. Under this correspondence between $\mathcal{G}_{1}$ and the class of unsigned graphs, the family $\mathcal{G}_{1}(\lambda)$ has a finite forbidden subgraph characterization within $\mathcal{G}_{1}$ if and only if the family of graphs with spectral radius at most $\lambda$ has a finite forbidden subgraph characterization. Therefore, the case where $p=1$ reduces to that in Theorem 5.
By restricting signed graphs in $\mathcal{G}_{2}$ to $\mathcal{G}_{1}$, we observe that if $\mathcal{G}_{2}(\lambda)$ has a finite forbidden subgraph characterization within $\mathcal{G}_{2}$, then so does $\mathcal{G}_{1}(\lambda)$ within $\mathcal{G}_{1}$. Conversely, suppose that $\mathcal{G}_{1}(\lambda)$ has a finite forbidden subgraph characterization $\mathcal{F}_{1}$ within $\mathcal{G}_{1}$. Let $\mathcal{F}_{2}$ be the family of signed graphs that are switching equivalent to signed graphs in $\mathcal{F}_{1}$. Using the fact that every $2$-colorable signed graph is switching equivalent to a $1$-colorable signed graph, one can check that $\mathcal{F}_{2}$ is a finite forbidden subgraph characterization of $\mathcal{G}_{2}(\lambda)$ within $\mathcal{G}_{2}$. ∎
The rest of the section focuses on solving 2 for $p\geq 3$.
Theorem 7 (Theorem 1.5 of Jiang and Polyanskii [6]).
The family $\mathcal{G}(\lambda)$ of signed graphs with smallest eigenvalue at least $-\lambda$ has a finite subgraph characterization (within the class of signed graphs) if and only if $\lambda<\lambda^{*}$. ∎
Proof of Theorem 3 for $\lambda<\lambda^{*}$.
In view of Theorem 7, we know that the family $\mathcal{G}(\lambda)$ has a finite forbidden subgraph characterization. By restricting signed graphs to those in $\mathcal{G}_{p}$, we obtain a finite forbidden subgraph characterization of $\mathcal{G}_{p}(\lambda)$ within $\mathcal{G}_{p}$ from that of $\mathcal{G}(\lambda)$. ∎
Proof of Theorem 3 for $\lambda\geq\lambda^{*}$ assuming Theorem 4.
Assume for the sake of contradiction that $\mathcal{F}$ is a finite forbidden subgraph characterization of $\mathcal{G}_{p}(\lambda)$ within $\mathcal{G}_{p}$. Because every signed graph in $\mathcal{G}_{p}\setminus\mathcal{G}_{p}(\lambda)$ contains a member of $\mathcal{F}$ as an induced subgraph, no $p$-colorable signed graph has its smallest eigenvalue in the open interval $(\max\left\{\lambda_{1}(F)\colon F\in\mathcal{F}\right\},-\lambda)$, which contradicts with Theorem 4. ∎
3 Binary rowing graphs
A large chunk of Theorem 4 is essentially established by Shearer [9], who proved that the set of spectral radii of all graphs is dense in $(\lambda^{\prime},\infty)$. As was pointed out in [3], Shearer actually proved that the set of spectral radii of all caterpillar trees [^1] A caterpillar tree is a tree in which all the vertices are within distance $1$ of its central path. is already dense in $(\lambda^{\prime},\infty)$. Since a caterpillar tree is bipartite (or equivalently, $2$-colorable), we rephrase Shearer’s result in terms of smallest eigenvalues.
Theorem 8 (Shearer [9]; cf. Theorem 3 of Doob [3]).
For every $\lambda\geq\lambda^{\prime}$ and every $\varepsilon>0$, there exists a caterpillar tree $G$ such that $\lambda_{1}(G)\in(-\lambda-\varepsilon,-\lambda)$. ∎
Jiang and Polyanskii introduced rowing graphs in [6] to show that the smallest eigenvalues of all graphs is dense in $(-\infty,-\lambda^{*})$, filling the gap between $-\lambda^{\prime}\approx-2.05817$ and $-\lambda^{*}\approx-2.01980$. We use the Kleene star of $\Sigma$, denoted $\Sigma^{*}$, to represent the set of all strings of finite length consisting of symbols in $\Sigma$, including the empty string.
Definition 9 (Rooted graphs and rowing graphs).
A rooted graph is a graph in which one vertex has been distinguished as the root. Given a rooted graph $F$, $n\in\mathbb{N}$, and a string $a\in\mathbb{N}^{n}$, a rowing graph $(F,a)$ is obtained from $F$ by attaching a path $v_{0}\dots v_{n}$ of length $n$ to the root $v_{0}$ of $F$, and attaching a clique of order $a_{i}$ to both $v_{i-1}$ and $v_{i}$ for every $i\in\left\{1,\dots,n\right\}$, and moreover the rowing graph $(F,a)$ is viewed as a rooted graph with root $v_{n}$.
Figure 1: The rooted graph $F_{1}$ and a schematic drawing of the rowing graph $(F_{1},20240125)$.
Under this convention, we may view the rowing graph $(F,ab)$ of $F$ as the rowing graph $((F,a),b)$ of the rowing graph $(F,a)$, for every $a,b\in\mathbb{N}^{*}$. See Figure 1 for an example of a rooted graph $F_{1}$ and a rowing graph of $F_{1}$.
To control chromatic numbers, we focus on a specific family of rowing graphs.
Definition 10 (Binary rowing graphs).
A binary rowing graph of a rooted graph $F$ is a rowing graph $(F,a)$ for which $a\in\left\{0,1\right\}^{*}$.
We import the following two ingredients from [6].
Lemma 11 (Lemma 2.22(c) of Jiang and Polyanskii [6]).
For every $\varepsilon>0$ there exists $\ell\in\mathbb{N}^{+}$ such that for every rooted graph $F$ and $a,b\in\mathbb{N}^{*}$,
$\lambda_{1}(F,a0^{\ell})<\lambda_{1}(F,a0^{\ell}b)+\varepsilon.\qed$
Lemma 12 (Lemma 2.22(e) of Jiang and Polyanskii [6]).
For every $\varepsilon>0$, there exists $m\in\mathbb{N}^{+}$ such that for every rooted graph $F$, every $n\geq m$, and every $a\in\mathbb{N}^{n}$, there exists $k\in\left\{1,\dots,m\right\}$ such that
$\lambda_{1}(F,a_{1}\dots a_{k-1}0a_{k}\dots a_{m})<\lambda_{1}(F,a)+\varepsilon.\qed$
Remark.
We point out that although these two lemmas were proved in [6] specifically for the rooted graph $F_{1}$ in Figure 1, the proofs over there actually work for arbitrary rooted graphs.
We adopt the following handy notation for every rooted graph $F$ and every $a\in\mathbb{N}^{*}$,
$\lambda_{1}(F,a0^{\infty}):=\lim_{n\to\infty}\lambda_{1}(F,a0^{n}).$
Note that the limit exists because $\lambda_{1}(F,a0^{n})$ decreases as $n$ increases, and the maximum degree of $(F,a0^{n})$ is at most that of $(F,a0^{2})$.
The last two ingredients can respectively be seen as variations of [6, Lemma 2.22(d)] and [6, Lemma 2.22(b)] adapted for binary rowing graphs.
Lemma 13.
For every rooted graph $F$ and every $a\in\left\{0,1\right\}^{*}$, if $-\lambda^{\prime}<\lambda_{1}(F,a01^{3}0^{\infty})<-2$, then
$\lambda_{1}(F,a01^{3}0^{\infty})\leq\lambda_{1}(F,a10^{\infty}).$
Lemma 14.
For the rooted graphs $F_{1},F_{2},F_{3}$ in Figure 2, the intervals
$\left(\lambda_{1}(F_{i},1^{8}),\lambda_{1}(F_{i},0^{\infty})\right)\text{ for }i\in\left\{1,2,3\right\}$
cover $(-\lambda^{\prime},-\lambda^{*})$. ∎
3124
21543
512346
Figure 2: Rooted graphs $F_{1},F_{2},F_{3}$.
We shall come back to the technical proofs of Lemmas 13 and 14 in the next section.
Proof of Theorem 4 assuming Lemmas 13 and 14.
In view of Theorem 8, we may assume that $\lambda\in(\lambda^{*},\lambda^{\prime})$. According to Lemma 14, there exists a rooted graph $F$ in Figure 2 such that $\lambda\in(-\lambda_{1}(F,0^{\infty}),-\lambda_{1}(F,1^{8}))$. Clearly, every binary rowing graph of $F$ is $3$-colorable.
Fix $\varepsilon>0$. We assume for the sake of contradiction that no binary rowing graph of $F$ has its smallest eigenvalue in $(-\lambda-\varepsilon,-\lambda)$. Let $\ell\in\mathbb{N}^{+}$ and $m\in\mathbb{N}^{+}$ be given by Lemmas 11 and 12. Without loss of generality, we may assume that $m+\ell\geq 8$ and $\ell\geq 3$. Define
$S=\left\{a_{1}\dots a_{m}b_{1}\dots b_{\ell}\in\left\{0,1\right\}^{m+\ell}\colon\lambda_{1}(F,a_{1}\dots a_{m}b_{1}\dots b_{\ell}c)<-\lambda\text{ for some }c\in\left\{0\right\}^{*}\right\}.$
Since $\lambda_{1}(F,1^{8})<-\lambda$, we must have $1^{m+\ell}\in S$, and so $S$ is non-empty. Let $ab$ be the minimum element of $S$ under the lexicographical order, where $a\in\left\{0,1\right\}^{m}$ and $b\in\left\{0,1\right\}^{\ell}$. Let $c\in\left\{0\right\}^{*}$ be the witness of $ab\in S$, that is, $\lambda_{1}(F,abc)<-\lambda$. Since no binary rowing graph of $F$ has its smallest eigenvalue in $(-\lambda-\varepsilon,-\lambda)$, we shall repeatedly use the fact that $\lambda_{1}(F,abc)+\varepsilon<-\lambda$.
We claim that $b=0^{\ell}$. Indeed, from Lemma 12 we obtain $k\in\left\{1,\dots,m\right\}$ such that
$\lambda_{1}(F,a_{1}\dots a_{k-1}0a_{k}\dots a_{m}bc)<\lambda_{1}(F,abc)+\varepsilon<-\lambda,$
and so $a_{1}\dots a_{k-1}0a_{k}\dots a_{m}b_{1}\dots b_{\ell-1}\in S$. By the minimality of $ab$ in $S$, we must have $a_{k}=\dots=a_{m}=b_{1}=\dots=b_{\ell}=0$.
In view of the claim, we get from Lemma 11 that
$\lambda_{1}(F,ab)<\lambda_{1}(F,abc)+\varepsilon<-\lambda.$ (1)
Since $\lambda_{1}(F,0^{m+\ell})>\lambda_{1}(F,0^{\infty})>-\lambda$, we must have $a\neq 0^{m}$. Let $k\in\left\{1,\dots,m\right\}$ be the last index of $a$ for which $a_{k}=1$. Set $\tilde{a}=a_{1}\dots a_{k-1}$, and set $n:=m+\ell-k$. Recall that $\ell\geq 3$, and so $n\geq 3$. By the minimality of $ab$, which is equal to $\tilde{a}10^{n}$, in $S$, we know that $\tilde{a}01^{3}0^{n-3}\in\left\{0,1\right\}^{m+\ell}\setminus S$, and, in particular, $\lambda_{1}(F,\tilde{a}01^{3}0^{\infty})\geq-\lambda$. Note that
$-\lambda^{\prime}<-\lambda\leq\lambda_{1}(F,\tilde{a}01^{3}0^{\infty})<\lambda_{1}(F,0^{\infty})<-2,$
We apply Lemma 13 to obtain that
$\lambda_{1}(F,\tilde{a}01^{3}0^{\infty})\leq\lambda_{1}(F,\tilde{a}10^{\infty})\leq\lambda_{1}(F,ab),$
which contradicts with (1). ∎
4 Proof of Lemmas 13 and 14
The following linear algebraic lemma characterizes $\lambda_{1}(F,a0^{\infty})$. Denote by $E_{v,v}$ the unit matrix where the $(v,v)$-entry with value $1$ is the only nonzero entry. We write $A\succeq 0$ when $A$ is positive-semidefinite and $A\succ 0$ when $A$ is positive-definite.
Lemma 15 (Lemma 26 of Acharya and Jiang [1]).
For every rooted graph $G$, and every $x\geq 2$,
$\lambda_{1}(G,0^{\infty})\geq-x\text{ if and only if }A_{G}+xI-yE_{v,v}\succeq 0,$
where $v$ is the root of $G$, and $y=x/2-\sqrt{x^{2}/4-1}$. ∎
Notice that each diagonal entry of $A_{G}+xI-yE_{v,v}$ increases as $x$ increases. We have the following version of Lemma 15 with strict inequalities.
Corollary 16.
For every rooted graph $G$, and every $x>2$,
$\lambda_{1}(G,0^{\infty})>-x\text{ if and only if }A_{G}+xI-yE_{v,v}\succ 0,$
where $v$ is the root of $G$, and $y=(x/2-\sqrt{x^{2}/4-1})$. ∎
We are in the position to characterize $\lambda_{1}(F,a01^{3}0^{\infty})$ and $\lambda_{1}(F,a10^{\infty})$ in Lemma 13.
Lemma 17.
For every rooted graph $G$, and every $x>2$,
$\lambda_{1}(G,01^{3}0^{\infty})\geq-x\text{ if and only if }A_{G}+xI-\alpha E_{v,v}\succeq 0,$
$\lambda_{1}(G,10^{\infty})\geq-x\text{ if and only if }A_{G}+xI-\beta E_{v,v}\succeq 0,$
where $v$ is the root of $G$,
$\begin{gathered}\alpha=\frac{(x^{5}-5x^{3}+2x^{2}+3x)y-(x^{6}-7x^{4}+4x^{3}+9x^{2}-6x-1)}{(x^{6}-7x^{4}+4x^{3}+9x^{2}-6x-1)y-(x^{7}-9x^{5}+6x^{4}+19x^{3}-20x^{2}-3x+6)},\\ \beta=\frac{y-(2x-2)}{xy-(x^{2}-1)},\text{ and }y=x/2-\sqrt{x^{2}/4-1}.\end{gathered}$ (2)
Proof.
Suppose that $x>2$. By Lemma 15, we know that
$\lambda_{1}(G,01^{3}0^{\infty})\geq-x\text{ if and only if }M_{1}:=A_{(G,01^{3})}+xI-yE_{u,u}\succeq 0,$
where $u$ is the root of $(G,01^{3})$. We partition the matrix $M_{1}$ into the following blocks
$M_{1}=\begin{pmatrix}A_{G}+xI&B_{1}\\ B_{1}^{\intercal}&C_{1}\end{pmatrix},\text{ where }C_{1}=A_{(\bullet,1^{3})}+xI-yE_{u,u}.$
Since $(\bullet,1^{3}0^{n})$ is a line graph for every $n\in\mathbb{N}$, we have $\lambda_{1}(\bullet,1^{3}0^{\infty})\geq-2>-x$, and so $C_{1}\succ 0$ via Corollary 16. Therefore $M_{1}\succeq 0$ if and only if the Schur complement $A_{G}+xI-B_{1}C_{1}^{-1}B_{1}^{\intercal}\succeq 0$. Let $v_{0}$ be the vertex in $V(G,0)\setminus V(G)$. Since the only nonzero entry of $B_{1}$ is its $(v,v_{0})$ entry, the matrix $B_{1}C_{1}^{-1}B_{1}^{\intercal}$ simplifies to $(C_{1}^{-1})_{v_{0},v_{0}}E_{v,v}$. Cramer’s rule yields $(C_{1}^{-1})_{v_{0},v_{0}}=\det C_{1}^{\prime}/\det C_{1}$, where $C_{1}^{\prime}$ is obtained from $C_{1}$ by removing the $v_{0}$-th row and column. A routine calculation yields
$\det C_{1}^{\prime}$ $=-(x^{5}-5x^{3}+2x^{2}+3x)y+(x^{6}-7x^{4}+4x^{3}+9x^{2}-6x-1),$
$\det C_{1}$ $=-(x^{6}-7x^{4}+4x^{3}+9x^{2}-6x-1)y+(x^{7}-9x^{5}+6x^{4}+19x^{3}-20x^{2}-3x+6),$
and so $(C_{1}^{-1})_{v_{0},v_{0}}=\alpha$.
By Lemma 15, we know that $\lambda_{1}(G,10^{\infty})\geq-x$ if and only if $M_{2}:=A_{(G,1)}+xI-yE_{v_{1},v_{1}}\succeq 0$, where $v_{1}$ is the root of $(G,1)$. We partition $M_{2}$ into blocks
$M_{2}=\begin{pmatrix}A_{G}+xI&B_{2}\\ B_{2}^{\intercal}&C_{2}\end{pmatrix},\text{ where }C_{2}=A_{(\bullet,0)}+xI-yE_{v_{1},v_{1}}.$
A similar argument shows that $M_{2}\succeq 0$ if and only if $A_{G}+xI-B_{2}C_{2}^{-1}B_{2}^{\intercal}\succeq 0$. Let $v_{2}$ be the other vertex in $V(G,1)\setminus V(G)$. Since the only nonzero entries of $B_{2}$ are its $(v,v_{1})$ and $(v,v_{2})$ entries, the matrix $B_{2}C_{2}^{-1}B_{2}^{\intercal}$ simplifies to $\left(\sum_{u_{1},u_{2}}(C_{2}^{-1})_{u_{1},u_{2}}\right)E_{v,v}$. A routine calculation yields $\sum_{u_{1},u_{2}}(C_{2}^{-1})_{u_{1},u_{2}}=\beta$. ∎
To prove Lemma 13, we simply compare $\alpha$ and $\beta$ defined in Lemma 17.
Proof of Lemma 13.
Let $G$ be the rooted graph $(F,a)$, and let $v$ be the root of $G$. For every $x>2$, we know through Lemma 17 that $\lambda_{1}(G,01^{3}0^{\infty})\geq-x$ if and only if $A_{G}+xI-\alpha E_{v,v}\succeq 0$, and $\lambda_{1}(G,10^{\infty})\geq-x$ if and only if $A_{G}+xI-\beta E_{v,v}\succeq 0$, where $\alpha$, $\beta$, and $y$ are defined as in (2).
Since $E_{v,v}\succeq 0$, it suffices to show that $\alpha\geq\beta$ for $x\in(2,\lambda^{\prime})$. Note that $x=y+1/y$, and $x\in(2,\lambda^{\prime})$ if and only if $y\in(\varphi^{-1/2},1)$, where $\varphi=(\sqrt{5}+1)/2$. Replacing $x$ by $y+1/y$ in $\alpha-\beta$, we obtain the rational function
$\frac{y(y-1)^{2}(y^{12}-2y^{11}+2y^{10}+2y^{9}-5y^{8}+6y^{7}-4y^{6}+2y^{5}+3y^{4}-6y^{3}+4y^{2}-1)}{(y^{2}-y+1)(-y^{10}+y^{9}+y^{8}-2y^{7}+y^{6}-y^{5}+y^{4}+2y^{3}-3y^{2}+y+1)}.$
It is routine to check via Sturm’s theorem that each polynomial factor in the rational function is positive on $(7/9,1)\approx(0.77778,1)$, which contains $(\varphi^{-1/2},1)\approx(0.78615,1)$. ∎
To prove Lemma 14, we gather the following computation about the smallest eigenvalues of binary rowing graphs of rooted graphs in Figure 2.
Proposition 18 (Hoffman [4]).
For every $n\in\mathbb{N}$, as $n\to\infty$, the smallest eigenvalue of the binary rowing graph $(F_{1},0^{n})$ of the rooted graph $F_{1}$ in Figure 2 decreases to $-\lambda^{*}$. ∎
Proposition 19.
The rowing graphs of the rooted graphs $F_{1},F_{2},F_{3}$ in Figure 2 satisfy the following properties:
$\left(\lambda_{1}(F_{1},1^{8}),\lambda_{1}(F_{1},0^{\infty})\right)$ $\supseteq[-\tfrac{136}{67},-\lambda^{*}),$
$\left(\lambda_{1}(F_{2},1^{7}),\lambda_{1}(F_{2},0^{\infty})\right)$ $\supseteq[-\tfrac{178}{87},-\tfrac{136}{67}],$
$\left(\lambda_{1}(F_{3},1^{4}),\lambda_{1}(F_{3},0^{\infty})\right)$ $\supseteq(-\lambda^{\prime},-\tfrac{178}{87}].$
5252-52- 523434-34- 346969-69- 691061061061061212-12- 128282-82- 8299996363636377-7- 74848-48- 4866663636363655-5- 52626-26- 2644441818181844-4- 41010-10- 1033333333
99995252525277-7- 73838-38- 3855552727272744-4- 41919-19- 1933331313131333-3- 377-7- 72222222214141414282828287171-71- 712828282814141414
55-5- 52424-24- 2444441515151533-3- 399-9- 9333333331717-17- 1766661212-12- 12181818182626-26- 2635353535
Figure 3: Test vectors $z_{1},z_{2},z_{3}$ on $(F_{1},1^{8}),(F_{2},1^{7}),(F_{3},1^{4})$
Proof.
We assign test vectors $z_{1},z_{2},z_{3}$ on $(F_{1},1^{8}),(F_{2},1^{7}),(F_{3},1^{4})$ as in Figure 3. The corresponding Rayleigh quotients are $-\tfrac{72334}{35635}$, $-\tfrac{4315}{2109}$, $-\tfrac{1875}{911}$, which are less than $-\tfrac{136}{67}$, $-\tfrac{178}{87}$, $-\lambda^{\prime}$ respectively.
According to Proposition 18, we know that $\lambda_{1}(F_{1},0^{\infty})=-\lambda^{*}$. Set $x_{2}=\tfrac{136}{67}$ and $x_{3}=\tfrac{178}{87}$. According to Corollary 16, to prove that $\lambda_{1}(F_{2},0^{\infty})>-x_{2}$ and $\lambda_{1}(F_{3},0^{\infty})>-x_{3}$, it suffices to show that
$A_{F_{2}}+x_{2}I-y_{2}E_{v_{2},v_{2}}\succ 0\text{ and }A_{F_{3}}+x_{3}I-y_{3}E_{v_{3},v_{3}}\succ 0,$
where $y_{2}=x_{2}/2-\sqrt{x_{2}^{2}/4-1}$, $y_{3}=x_{3}/2-\sqrt{x_{3}^{2}/4-1}$, and $v_{2}$ and $v_{3}$ are the roots of $F_{2}$ and $F_{3}$ respectively. Since $y_{2}\approx 0.84151$, $y_{3}\approx 0.80734$, $E_{v_{2},v_{2}}\succeq 0$ and $E_{v_{3},v_{3}}\succeq 0$, it suffices to show that
$M_{2}:=A_{F_{2}}+x_{2}I-y_{2}^{\prime}E_{v_{2},v_{2}}\succ 0\text{ and }M_{3}:=A_{F_{3}}+x_{3}I-y_{3}^{\prime}E_{v_{3},v_{3}}\succ 0,$
where $y_{2}^{\prime}=\tfrac{85}{101}\approx 0.84158$ and $y_{3}^{\prime}=\tfrac{285}{353}\approx 0.80737$. Note that
$M_{2}=\begin{pmatrix}\tfrac{136}{67}&1&&&1\\ 1&\tfrac{136}{67}&1&&1\\ &1&\tfrac{136}{67}&1&1\\ &&1&\tfrac{136}{67}&1\\ 1&1&1&1&\tfrac{8041}{6767}\end{pmatrix}\text{ and }M_{3}=\begin{pmatrix}\tfrac{178}{87}&1&&&&\\ 1&\tfrac{178}{87}&1&&&\\ &1&\tfrac{178}{87}&1&&\\ &&1&\tfrac{178}{87}&&1\\ &&&&\tfrac{178}{87}&1\\ &&&1&1&\frac{38039}{30711}\end{pmatrix}.$
Notice that the leading principal submatrix of $M_{2}$ of order $4$ and that of $M_{3}$ of order $5$ are positive definite, because both submatrices are of the form $A_{G}+xI$, where $G$ is a graph of maximum degree $2$, and $x>2$. Finally, it is routine to compute that $\det M_{2}=\tfrac{50854155}{136362635807}>0$ and $\det M_{3}=\tfrac{23578825817}{153070048956177}>0$. ∎
Proof of Lemma 14.
It follows from Proposition 19 immediately. ∎
5 Spherical two-distance sets
A spherical two-distance set in $\mathbb{R}^{d}$ is a collection of unit vectors whose pairwise inner products lie in $\left\{\alpha,\beta\right\}$. There was some partial success specifically in the regime where $-1\leq\beta<0\leq\alpha<1$. For fixed $\alpha$ and $\beta$, denote by $N_{\alpha,\beta}(d)$ the maximum size of such a set in $\mathbb{R}^{d}$.
When $\beta=-\alpha$, the problem specializes to the famous question of equiangular lines with fixed angles, whose high-dimensional behavior is now understood thanks to the breakthrough of Jiang, Tidor, Yao, Zhang and Zhao [7]. Outside the equiangular setting, the landscape is much richer and remain largely open.
Recently, the limit $\lim_{d\to\infty}N_{\alpha,\beta}(d)/d$ was determined in [6, 8] when $p\leq 2$ or $\lambda<\lambda^{*}$, where the parameters are defined by
$p=\left\lfloor{-\frac{\alpha}{\beta}}\right\rfloor+1\quad\text{and}\quad\lambda=\frac{1-\alpha}{\alpha-\beta}.$
In [8], a spherical two-distance set with prescribed inner products $\alpha$ and $\beta$ in $\mathbb{R}^{d}$ was constructed, and it was conjectured to provide the optimal $N_{\alpha,\beta}(d)$ up to an additive error of $o(d)$. The limit $\lim_{d\to\infty}N_{\alpha,\beta}(d)/d$ is believed to depend on the following spectral graph theoretic quantity:
$k_{p}(\lambda)=\inf\left\{\frac{\lvert G\rvert}{\mathrm{mult}(-\lambda,G)}\colon G\in\mathcal{G}_{p}(\lambda)\right\},$
where $\mathrm{mult}(-\lambda,G)$ denotes the multiplicity of $-\lambda$ as an eigenvalue of $G$.
Conjecture 20 (Conjecture 1.11 of Jiang et al. [8]).
Fix $-1\leq\beta<0\leq\alpha<1$. Set $\lambda=(1-\alpha)/(\alpha-\beta)$ and $p=\lfloor-\alpha/\beta\rfloor+1$. Then
$N_{\alpha,\beta}(d)=\begin{cases}\dfrac{k_{p}(\lambda)d}{k_{p}(\lambda)-1}+o(d)&\text{if }k_{p}(\lambda)<\infty,\\ d+o(d)&\text{otherwise}.\end{cases}$
To prove 20, a forbidden-subgraph approach was discovered in [8] that reduces the estimation of $N_{\alpha,\beta}(d)$ to that of the following quantity. See [8, Section 5] for the reduction.
Definition 21 (Definition 5.2 of Jiang et al. [8]).
Given $p\in\mathbb{N}^{+}$ and a family $\mathcal{F}$ of signed graphs, let $M_{p,\mathcal{F}}(\lambda,N)$ be the maximum possible value of $\mathrm{mult}(-\lambda,G)$ over all $p$-colorable signed graphs $G$ on at most $N$ vertices that do not contain any member of $\mathcal{F}$ as an induced subgraph and satisfy $\lambda_{p+1}(G)\geq-\lambda$. Here $\lambda_{p+1}(G)$ denotes the $(p+1)$-th smallest eigenvalue of $G$.
Conjecture 22 (Conjecture 5.4 of Jiang et al. [8]).
For every $\lambda>0$ and $p\in\mathbb{N}^{+}$, there exists a finite family $\mathcal{F}$ of signed graphs with $\lambda_{1}(F)<-\lambda$ for each $F\in\mathcal{F}$ such that
$M_{p,\mathcal{F}}(\lambda,N)\leq\begin{cases}N/k_{p}(\lambda)+o(N)&\text{if }k_{p}(\lambda)<\infty,\\ o(N)&\text{otherwise}.\end{cases}$
When $\lambda<\lambda^{*}$, according to Theorem 7, the family $\mathcal{F}$ in [6] was taken as a finite forbidden subgraph characterization of the family of signed graphs with smallest eigenvalue at least $-\lambda$ (within the class of signed graphs). In view of Definition 21, for this specific $\mathcal{F}$, the quantity simplifies to
$M_{p,\mathcal{F}}(\lambda,N)=\max\left\{\mathrm{mult}(-\lambda,G)\colon\lvert G\rvert\leq N\text{ and }G\in\mathcal{G}_{p}(\lambda)\right\},$
which in turn confirms 22 through a simple argument. We refer the readers to [6, Section 4] for more details.
Since a $p$-colorable graph $G$ never contains a subgraph $F$ that is not $p$-colorable, it is more economical to choose $\mathcal{F}$ to be a forbidden subgraph characterization of $\mathcal{G}_{p}(\lambda)$ within $\mathcal{G}_{p}$. This is the motivation behind 2. Our Theorem 4 therefore shows that restricting to $p$-colorable signed graphs yields no new finite forbidden subgraph characterizations once $p\geq 3$.
Acknowledgements
We thank Xiaonan Liu for valuable early-stage discussions and gratefully acknowledge travel support from LSU Provost’s Fund for Innovation in Research—Seminar/Collaborator Support Fund.